Circular No. 6329

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) In adopting, originally on IAUC 6304, the magnitude formula (5.5 +

5 log Δ + 10 log r) for this comet, we were well aware that CCD observations were then running up to 2 mag fainter and visual estimates up to 1 mag brighter than this. This is still the case (see MPC 26606–26608, IAUC 6328), which mainly serves to confirm that the choice of the term 10 log r (r^{−4} variation) was not unreasonable. The following general orbit, from MPC 26724 and utilizing observations through Feb. 27, shows a 5σ departure from a parabolic solution, some worry about this determination arising from the use of the isolated Jan. 1 prediscovery observations:

Epoch = 1996 Apr. 27.0 TT

T = 1996 May 1.4025 TT  ω = 130.^{o}1925
e = 0.999846  Ω = 188.0514 }2000.0
q = 0.230123 AU  i = 124.9016

Nevertheless, the adjustment necessary to obtain the “original” barycentric orbit is quite large for this comet. The indications are thus that this is not a “new” comet in the Oort sense, and that it was last near the sun 10 000–20 000 years ago. This may inspire further confidence in the r^{−4} law, and since visual observations would seem more relevant than CCD observations in discussions of naked-eye visibility, one might expect the comet to become even brighter than predicted. It is important to note, however, that when the comet is closest to the earth (0.102 AU on Mar. 25.3 UT), that brightness is likely to be diffused over an area several degrees across. More typical tail development is likely to come later as the comet approaches the sun. The comet is unique in its close passage to the earth and subsequent small perihelion distance. If the brightness holds, it will be the intrinsically brightest comet to pass so close to the earth since 1556. The ephemeris uncertainty is still at least 1′.

1996 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 13 14^{h}55.^{m}36 −16^{o}20.^{′}4 0.422 1.290 126.^{o}6 38.^{o}2 4.7
14 14 55.40 −15 10.7 0.389 1.271 127.9 38.1 4.5
15 14 55.38 −13 48.0 0.357 1.251 129.2 38.0 4.2
16 14 55.26 −12 08.8 0.325 1.232 130.6 37.8 4.0
17 14 55.04 −10 07.9 0.294 1.212 132.1 37.5 3.7
18 14 54.69 − 7 38.0 0.263 1.192 133.6 37.2 3.4
19 14 54.16 − 4 28.2 0.232 1.172 135.1 36.9 3.0
20 14 53.38 − 0 22.4 0.203 1.152 136.4 36.6 2.7
1996 February 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden